in a region of space has constant charge density, p. Assume that the region is infinite in the x-y plane and of height, d. Determine the electric field inside and outside the charged region.


1
Expert's answer
2021-10-24T18:26:10-0400

Inside.

"\oint EdA=\frac {q_{in}}{\epsilon _0},"

"dq=\rho dV=\rho xl^2dx,"

"q=\int dq=\frac{\rho l^2d^2}8,"

"E=\frac{\rho d^2}{8\epsilon _0}."

Outside.

"\oint EdA=\frac{q_{in}}{\epsilon_0},"

"q=\int dq=\frac{\rho l^2a^2}{2},"

"E=\frac{\rho a^2}{2\epsilon_0}."